Solve by completing the square
x2 + 6x – 4 =0

Answer by richwmiller(17219)

x^2+6x=4
1/2*6=3
3^2=9
x^2+6x+9=13
(x+3)^2=13
x+3=|sqrt(13)|
x=-sqrt(13)-3
x=sqrt(13)-3